at what rate would it be necessary for a single conductor to cut the flux in order that a current of 1.2ma flows through it when 10 ohms resistor is connected.
a) 1.2 * 10^-3
​b) 1.2 * 10^-2
​c) 1.3 * 10^-3
d) 1.1 * 10^-3
 

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
e  = N d(phi)/dt
 
e = IR
 
d (phi)dt = e/N = 1.2 * 10^-2 wb
 
Hope it helps
 
Regards
Arun (askIITians forum expert)